OpenStudy (chrisplusian):

I am trying to show how (dV/dt)=v... where v=velocity, and (dV/dt) is the change in volume with respect to time.... PLEASE HELP!!!!

OpenStudy (anonymous):

its impossible to prove that u know why .. cause its wrong xD dV/dt = m^3/second and v = vleocity = m/second at the most dV/dt = vA , where v is the velocity of some fluid (maybe) and A is the area through which it is flowing!

OpenStudy (chrisplusian):

This is actually my exact problem... I am doing a physics project of fluid flow through a conical tank.... I have arrived attwo seperate equations.... one for torrcellis law and and one for the volume of a conical tank.... now I need to be able to tie them together to solve a differential equation and I am stuck

OpenStudy (mrnood):

Torricellis law gives you the velocity (v) through an orifice related to the depth of the tank and says that v= sqrt(2gh) The volume flow rate is therefore vA where A is area of the orifice. so if you have an expression for volume (V) in terms of h (for the conical tank) then you can work out dV/dt = A*sqrt(2gh)